### v1

# Dilaton-induced variations in Planck constant and speed of light

Abstract
We highlight a hidden aspect of scale-invariant actions that allow matter to couple with a dilaton field:   the dynamics of the dilaton can be used to induce variations in Planck’s quantum of action $\hbar$ and the speed of light $c$. This mechanism for generating variable $\hbar$ and $c$ in curved spacetimes offers novel insights into the origin of late-time cosmic acceleration, bypassing the need for dark energy.
Motivation—Let us revisit the quantum electrodynamics (QED) of a spinor field $\psi$ with inertial mass $m$ and charge $e$, coupled with a $U(1)$ gauge vector field $A_{\mu}$ and embedded in a 4-dimensional spacetime described by the Einstein–Hilbert (EH) action of General Relativity (GR). With the participation of $\hbar$ and $c$ explicitly restored, the full action reads:
$\mathcal{S}_{1}$ $=\int d^{4}x\sqrt{-g}\,\Bigl{[}\mathcal{L}_{\text{EH}}+\mathcal{L}_{\text{QED}}\Bigr{]}$ (1)
$\mathcal{L}_{\text{EH}}$ $=\frac{c^{3}}{16\pi\hbar G}\,\mathcal{R}$ (2)
$\mathcal{L}_{\text{QED}}$ $=i\,\bar{\psi}\gamma^{\mu}\nabla_{\mu}\psi+\frac{e}{\sqrt{\hbar c}}\bar{\psi}\gamma^{\mu}A_{\mu}\psi+m\frac{c}{\hbar}\,\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ (3)
$F_{\mu\nu}$ $:=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ (4)
The restoration of $\hbar$ and $c$ in these expressions will be instrumental for our later discussion when considering modified theories of gravity and matter. For instance, in place of $\mathcal{L}_{\text{EH}}$, Brans and Dicke proposed an extension of GR known as the Brans–Dicke (BD) gravity action [1]:
$\mathcal{L}_{\text{BD}}=\phi\,\mathcal{R}-\frac{\omega}{\phi}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$ (5)
Traditionally, the dynamical BD scalar field $\phi$ is regarded as an inverse (variable) Newton gravitational constant $G$. However, by comparing the right-hand-side of Eq. (2) with the first term in the right-hand-side of Eq. (5), it becomes evident that the combination $c^{3}/(\hbar G)$, and not merely $1/G$, should be associated with $\phi$. If the matter Lagrangian only allows matter to couple minimally with gravity—such as $\mathcal{L}_{\text{QED}}$ in Eq. (3)—the inertial mass $m$ and charge $e$, along with $\hbar$ and $c$, are unrelated to $\phi$; in this situation, $1/G$ can be identified with $\phi$, as initially proposed by Brans and Dicke.
However, this relationship between $G$ and $\phi$ could cease to hold if matter couples non-minimally with gravity via the BD field. In this scenario, the involvement of $\phi$ in the matter Lagrangian, in principle, can cause both $c$ and $\hbar$ to vary alongside $\phi$ in spacetime, representing a departure from the standard BD paradigm. The current paper explores this potentially far-reaching scenario.
Derivation—For convenience, let us replace the BD field $\phi$ with a scalar field $\chi$ via the substitution $\phi:=\chi^{2}$, transforming the BD action into
$\mathcal{L}_{\chi}=\chi^{2}\,\mathcal{R}-4\omega\,g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi$ (6)
Furthermore, we allow the quadratic term $\bar{\psi}\psi$ to couple non-minimally with gravity via the field $\chi$ in the form $\chi\,\bar{\psi}\psi$. The modified QED Lagrangian is given by
$\mathcal{L}_{\text{QED}}^{(\chi)}=i\,\bar{\psi}\gamma^{\mu}\nabla_{\mu}\psi+\sqrt{\alpha}\,\bar{\psi}\gamma^{\mu}A_{\mu}\psi+\mu\,\chi\,\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ (7)
where the Dirac gamma matrices satisfy $\gamma^{\mu}\gamma^{\nu}+\gamma^{\nu}\gamma^{\mu}=2\,g^{\mu\nu}$, and the spacetime covariant derivative $\nabla_{\mu}$ acts on the spinor via vierbein and spin connection. The full action now reads
$\mathcal{S}_{2}$ $=\int d^{4}x\sqrt{-g}\,\Bigl{[}\mathcal{L}_{\chi}+\mathcal{L}_{\text{QED}}^{(\chi)}\Bigr{]}$ (8)
All parameters $\omega$, $\alpha$, and $\mu$ in $\mathcal{L}_{\chi}$ and $\mathcal{L}_{\text{QED}}^{(\chi)}$ are dimensionless. The modified action $\mathcal{S}_{2}$ is scale invariant; i.e., it remains unchanged under a global Weyl rescaling:
$g_{\mu\nu}\rightarrow\Omega^{2}g_{\mu\nu};\ \psi\rightarrow\Omega^{-3/2}\psi;\ A_{\mu}\rightarrow A_{\mu};\ \chi\rightarrow\Omega^{-1}\chi$ (9)
In the literature, when scale symmetry, also known as dilatation symmetry, is broken—specifically, when $\chi$ spontaneously acquires a non-vanishing vacuum expectation value—the process gives rise to a Goldstone mode conventionally referred to as a dilaton. We will adopt the terminology of “dilaton” for $\chi$ in this paper. We should note, however, that the field $\chi$ in our action $\mathcal{S}_{2}$ can vary in spacetime, hence forfeiting its translational invariance.
Interestingly, a scale-invariant action of gravity and matter, such as $\mathcal{S}_{2}$, is able to evade observational bounds on the fifth force, a result established in [2, 3]. One can also add a “potential” term $V(\chi)$ to $\mathcal{L}_{\chi}$ provided that the added term respects scale symmetry. For example, $V(\chi)$ may contain terms like $\chi^{4}$ and $\chi^{-4}(\nabla\chi)^{4}$. Nevertheless, the inclusion of a scale-invariant potential will not affect the consideration presented in the rest of this paper.
Identification—Consider the dilaton $\chi$ as a slowly varying background field in spacetime. A comparison of $\mathcal{L}_{\chi}$ versus $\mathcal{L}_{\text{EH}}$ and $\mathcal{L}_{\text{QED}}^{(\chi)}$ versus $\mathcal{L}_{\text{QED}}$ results in the following identification:
$\frac{c^{3}}{16\pi\hbar G}:=\chi^{2};\ \ \ \frac{e}{\sqrt{\hbar c}}:=\sqrt{\alpha};\ \ \ m\frac{c}{\hbar}:=\mu\,\chi$ (10)
We will present two alternative routes to meet these identities.
The Fujii–Wetterich scheme: Under the canonical assumption that $\hbar$ and $c$ are constants, one then obtains from (10):
$e=\bigl{(}\alpha\hbar c\bigr{)}^{1/2};\ \ \ m_{\chi}=\frac{\mu\hbar}{c}\,\chi$ (11)
and
$G_{\chi}=\frac{c^{3}}{16\pi\hbar}\,\chi^{-2}\,.$ (12)
Here, the subscript $\chi$ in $m_{\chi}$ and $G_{\chi}$ underlines the dependence of $m$ and $G$ on $\chi$. Notably, the gauge charge $e$ remains unrelated to $\chi$. To the best of our knowledge, this scheme, which results in variable $G$ and variable mass, was first laid out by Fujii [4] and Wetterich [5, 6, 7, 8, 9] although they approached it using matter–gravity actions different from $\mathcal{S}_{2}$. The Planck mass, defined as $m_{\chi}^{\text{Planck}}\,:=$$\sqrt{\frac{\hbar c}{G_{\chi}}}=$$\,\sqrt{16\pi}\frac{\hbar}{c}\,\chi$, is variable in this Fujii–Wetterich (FW) scheme.
It is important to note that in this scheme, owing to the relationship involving $m_{\chi}$ in Eq. (11), the dilaton can only affect massive particles, viz. the quanta of the spinor field $\psi$. The dilaton does not impact massless particles, viz. the quanta of the $U(1)$ gauge vector field $A_{\mu}$.
Our scheme: There is no a priori theoretical reason to prevent Planck’s quantum of action and the speed of light from associating with the dilaton, however. Instead, it is permissible to relate $\hbar$ and $c$ to $\chi$ via the following assignments (with the subscript $\chi$ attached):
$\hbar_{\chi}:=\hat{\hbar}\ \Bigl{(}\frac{\chi}{\hat{\chi}}\Bigr{)}^{-1/2};\ \ \ c_{\chi}:=\hat{c}\ \Bigl{(}\frac{\chi}{\hat{\chi}}\Bigr{)}^{1/2}$ (13)
where $\hat{\hbar}$ and $\hat{c}$ represent the values of $\hbar_{\chi}$ and $c_{\chi}$ at a reference point where $\chi=\hat{\chi}$ (with $\hat{\chi}\neq 0$). Employing Eq. (10), our assignments unambiguously lead to
$e=\bigl{(}\alpha\hat{\hbar}\hat{c}\bigr{)}^{1/2};\ \ \ m=\frac{\mu\,\hat{\chi}\hat{\hbar}}{\hat{c}}$ (14)
and
$G=\frac{\hat{c}^{3}}{16\pi\hat{\chi}^{2}\hat{\hbar}}\,.$ (15)
Importantly, per Eq. (14), the charge $e$ and inertial mass $m$ are independent of $\chi$. This result is desirable, as both charge and inertial mass—being intrinsic properties of a particle—should be oblivious to external factors, in particular the background dilaton field $\chi$. Befittingly, in our scheme, rather than making mass variable, the effect of the dilaton is translated into $\hbar_{\chi}$ and $c_{\chi}$, which respectively regulate the quantization and propagation of fields against a background spacetime. [Note: It is essential to note that the dimensionless parameters $\alpha$ and $\mu$ are “running coupling constants” in the renormalization group flow of $\mathcal{L}_{\text{QED}}^{(\chi)}$ when loop corrections involving $\psi$ and $A_{\mu}$ are included. Thus, although $e$ and $m$ are independent of $\chi$, they can “run” as functions of the momentum scale at which they are measured.]
The role of $\chi$ in determining $\hbar$ and $c$ underscores a stark distinction between the two schemes. As mentioned earlier, in the FW scheme, the dilaton—by design—can only affect massive particles while leaving massless particles unaffected. In contrast, our scheme allows the dilaton—through its involvement with $\hbar$ and $c$—to govern the quantization and propagation of all fields, irrespective of their mass parameters. The ability to make the dilaton universally affect all types of fields is a major virtue of our scheme in comparison to the FW scheme.
Moreover, our scheme brings about an additional benefit: not only is the $U(1)$ gauge charge $e$—which specifies the strength of electromagnetic interaction—independent of $\chi$, but the Newton gravitational constant $G$—characterizing the strength of gravitational interaction—is also independent of $\chi$, as evident in Eq. (15). Note that in our scheme, the Planck mass, given by $m_{\chi}^{\text{Planck}}\,:=$$\sqrt{\frac{\hbar_{\chi}c_{\chi}}{G}}=$$\,\sqrt{16\pi}\,\frac{\hat{\chi}\hat{\hbar}}{\hat{c}}$, is constant rather than variable.  [^1] Our paper does not concern with the hierarchy issue, namely why $m_{\chi}^{\text{Planck}}\gg m$, or equivalently why $\mu\ll\sqrt{16\pi}$.
A mechanism to generate variable $\hbar$ and $c$—Consider an open set surrounding a given point $x^{*}$ on the spacetime manifold. With the dilaton $\chi$ varying sufficiently slowly, it can be treated as having a constant value within the open set. Utilizing Eq. (13) from our scheme, the Planck ‘constant’ and the speed of light acquire their respective values $h_{\chi}$ and $c_{\chi}$, to be applicable solely for this region. In a tangent frame to the manifold at $x^{*}$, our scheme maps $\mathcal{L}_{\text{QED}}^{(\chi)}$ to an effective Lagrangian given by
$\mathcal{L}_{\text{QED}}^{\text{ eff}}$ $=i\,\bar{\psi}\gamma^{\mu}\nabla_{\mu}\psi+\frac{e}{\sqrt{\hbar_{\chi}c_{\chi}}}\bar{\psi}\gamma^{\mu}A_{\mu}\psi+m\frac{c_{\chi}}{\hbar_{\chi}}\,\bar{\psi}\psi$
$-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ (16)
Within the open set, $\mathcal{L}_{\text{QED}}^{\text{ eff}}$ describes the quantum electrodynamics of a spinor field with charge $e$ and inertial mass $m$, specified in Eq. (14), coupled with a $U(1)$ gauge vector field. It is important to emphasize that, unlike the original $\mathcal{L}_{\text{QED}}$ in Eq. (3), the quantization and propagation of fields in $\mathcal{L}_{\text{QED}}^{\text{ eff}}$ are governed by the effective quantum of action $\hbar_{\chi}$ and speed of light $c_{\chi}$, which in turn are determined by the dilaton $\chi$ via Eq. (13). More generally, as $\chi$ varies across the manifold, different locations then correspond to separate replicas of the effective Lagrangian $\mathcal{L}_{\text{QED}}^{\text{ eff}}$, each operating with its respective values of $\hbar_{\chi}$ and $c_{\chi}$. Conceptually, the dynamics of $\chi$ thus induces variations in $\hbar$ and $c$ across spacetime.
The operational meaning of variable $c$ and $\hbar$—In 1911 Einstein originated the idea of a variable speed of light (VSL) during his formulation of GR [10, 11, 12]. Although he did not pursue this idea further, the possibility of VSL was revived by Dicke in 1957 [13], as well as by Moffat, Albrecht, and Magueijo in the 1990s [14, 15]. Importantly, as Einstein recognized in [11, 12], the VSL concept does not contradict the principle of the constancy of $c$ with respect to local Lorentz boosts;   this is because the Lorentz symmetry is only required to be valid locally, while the value of $c$ may vary from one point to another in spacetime. Furthermore, it is incorrect to dismiss the distinction between keeping $c$ constant or allowing it to vary as merely a trivial redefinition of units (i.e., rods and clocks): a varying $c$ leads to a refraction effect—a physically measurable phenomenon—that impacts the propagation of light rays, whereas a constant $c$ does not.
A variable $\hbar$—as a function of $\chi$—would affect the quantization of fields, specifically the commutation relation of position and momentum for particles, given by
$\hat{x}\,\hat{p}-\hat{p}\,\hat{x}=i\,\hbar_{\chi}\,.$
In relation to this, a variable $\hbar$ would influence the time evolution of a quantum state $\bigl{|}\Psi\bigr{\rangle}$ according to
$i\hbar_{\chi}\,\frac{\partial}{\partial t}\bigl{|}\Psi\bigr{\rangle}=\hat{H}\bigl{|}\Psi\bigr{\rangle}\,.$ (17)
A physical consequence is that, for two locations with different dilaton values $\chi_{1}$ and $\chi_{2}$, the evolution of a quantum state occurs at different rates determined by each respective $h_{\chi}$ value. The discrepancy in the clock rates at the two locations gives rise to a new time dilation effect, resulting from the variation in $\hbar$ induced by the dynamics of $\chi$. We will briefly discuss this concrete definitive prediction in a later section of the paper.
Illustration using the Hydrogen atom—By applying the variational principle to the spinor field and the gauge vector field in the effective $\mathcal{L}_{\text{QED}}^{\text{ eff}}$ given in Eq. (16), it is straightforward to derive the Dirac equation
$\Biggl{(}i\gamma^{\mu}\partial_{\mu}+\frac{e}{\sqrt{\hbar_{\chi}c_{\chi}}}\gamma^{\mu}A_{\mu}+m\,\frac{c_{\chi}}{\hbar_{\chi}}\Biggr{)}\,\psi$ $=0$ (18)
and the Maxwell equation
$\partial_{\nu}F^{\nu\mu}=\frac{e}{\sqrt{\hbar_{\chi}c_{\chi}}}\,\,\bar{\psi}\gamma^{\mu}\psi\,.$ (19)
For convenience, we will refer to $\psi$ as an electron field and $A_{\mu}$ as an electromagnetic (EM) field. For a Coulomb potential, where $A_{0}=-\frac{1}{\sqrt{\hbar_{\chi}c_{\chi}}}\frac{e}{r}$, $A_{1}=A_{2}=A_{3}=0$, the Dirac equation becomes
$\Biggl{(}i\gamma^{\mu}\partial_{\mu}-\gamma^{0}\frac{1}{\hbar_{\chi}c_{\chi}}\frac{e^{2}}{r}+m\,\frac{c_{\chi}}{\hbar_{\chi}}\Biggr{)}\,\psi=0$ (20)
which describes the motion of an electron in a hydrogen atom. The Bohr radius of its groundstate is given by
$a_{B}$ $=\frac{\hbar_{\chi}}{\alpha\,mc_{\chi}}=\frac{1}{\alpha\,\mu\,\chi}\propto\chi^{-1}$ (21)
which is reciprocal to $\chi$. More generally, it can be anticipated—from the ground of dimensionality—that the lengthscale $l$ of any physical process occurring within the open set of a constant $\chi$ is also reciprocal to $\chi$, viz.
$l\propto\chi^{-1}$ (22)
The energy level of an (relativistic) electron in a quantum state $\left|n,\,j\right\rangle$ of a Hydrogen atom is a well-established result; it is given by [16]
$E_{n}^{j}$ $=\mathcal{N}_{n}^{j}\,mc_{\chi}^{2}=\mathcal{N}_{n}^{j}\,\mu\hat{\hbar}\hat{c}\,\chi\propto\chi$ (23)
which is proportional to $\chi$. In the above expression, the factor $\mathcal{N}_{n}^{j}\!$ is equal to ${\scriptstyle{\scriptstyle\left(1+\alpha^{2}\left(n-j-\frac{1}{2}+\sqrt{\left(j+\frac{1}{2}\right)^{2}-\alpha^{2}}\right)^{-2}\right)^{-1/2}}}$ with $n=1,2,3,...$ and $j=\frac{1}{2},\frac{3}{2},\frac{5}{2},...$.
Anisotropic scaling in the clock rate—Induced by electric dipole, a hydrogen atom in the excited state $\left|n=2,\,j=3/2\right\rangle$ can spontaneously transition to the groundstate $\left|n=1,\,j=1/2\right\rangle$, as allowed by the selection rule $\Delta j=\pm 1$. The energy of the photon emitted is $\Delta E=E_{n=2}^{j=3/2}-E_{n=1}^{j=1/2}$, and the frequency of the emitted photon is
$\nu$ $=\frac{\Delta E}{\hbar_{\chi}}=\Bigl{(}\mathcal{N}_{n=2}^{j=3/2}-\mathcal{N}_{n=1}^{j=1/2}\Bigr{)}\,\frac{\mu\hat{c}}{\hat{\chi}^{1/2}}\,\chi^{3/2}\propto\chi^{3/2}$ (24)
Therefore, the propagation of the photon within the open set of constant $\chi$ has a timescale that behaves as
$\tau:=\frac{1}{\nu}\propto\chi^{-3/2}$ (25)
Similarly, the rate of transition (i.e. Einstein’s coefficient) from the initial state $\left|i\right\rangle=\left|n=2,\,j=3/2\right\rangle$ to the final state $\left|f\right\rangle=\left|n=1,\,j=1/2\right\rangle$ can be found to be [17]:
$A=\frac{4\alpha}{3}\frac{m^{3}c_{\chi}^{4}}{h_{\chi}^{3}}\left\langle\vec{r}_{if}\right\rangle^{2}\,.$ (26)
where the matrix element of the electric dipole is $\left\langle\vec{r}_{if}\right\rangle:=\left\langle f|\vec{r}|i\right\rangle$. Given that $\left\langle\vec{r}_{if}\right\rangle\propto\chi^{-1}$, $c_{\chi}\propto\chi^{1/2}$, and $h_{\chi}\propto\chi^{-1/2}$, the transition rate behaves as $A\propto\chi^{3/2}$. Thus, the half-life of the decay process for an unstable quantum system scales as
$\tau:=\frac{1}{A}\propto\chi^{-3/2}$ (27)
in alignment with Eq. (25). Although this result was derived through illustrations involving a hydrogen atom, the time scaling behavior is generic owing to the time evolution of quantum states, given in Eq. (17). With the Hamiltonian scales similarly to energy (i.e., $\hat{H}\propto\chi$ per Eq. (23)), and $\hbar_{\chi}\propto\chi^{-1/2}$, the timescale of the evolution of a quantum state is thus $\tau\propto\hat{H}/\hbar_{\chi}\propto\chi^{3/2}$.
Moreover, this time scaling law can also be understood from the perspective of variable $c$. The timescale of a physical process in the open set of a constant $\chi$ is deducible from its lengthscale $l$ (with $l\propto\chi^{-1}$) and the effective speed of light $c_{\chi}$. That is to say
$\tau:=\frac{l}{c_{\chi}}\propto\chi^{-3/2}\,.$ (28)
in perfect agreement with Eq. (25).
Finally, combining Eq. (28) with $l\propto\chi^{-1}$, we arrive at an anisotropic relationship between the timescale $\tau$ and lengthscale $l$ of a given physical process
[…]
This result is one of the key findings of our paper.  [^2] There are two examples in classical physics that carry a $3/2$–exponent hallmark: (1) Kepler’s Third law, which states that the square of a planet’s orbital period $T$ is proportional to the cube of the semi-major axis $a$ of its orbit, viz. $T^{2}\propto a^{3}$, or $T\propto a^{3/2}$.(2) The cosmic factor $a$ of a spatially flat Einstein–de Sitter universe evolves as $a\propto t^{2/3}$, or $t\propto a^{3/2}$. These behaviors resemble the anisotropic time scaling, $\tau\propto l^{3/2}$, discussed in this paper.
Prediction: A new time dilation effect (of the Third kind)—The time scaling law (29) indicates that the rate of any clock—be it mechanical, electronic, or atomic—varies in spacetime, as a function of $\chi$, in an anisotropic fashion in comparison to the length of a rod. Mathematically, while the length of a rod scales as $l\propto\chi^{-1}$, the rate of a clock scales as $\tau^{-1}\propto\chi^{3/2}$, with the $3/2$–exponent arising due to $c_{\chi}\propto\chi^{1/2}$ (or, equivalently, due to $\hbar_{\chi}\propto\chi^{-1/2}$).
In principle, this effect can be measured experimentally:  Prepare two identical clocks at a location $A$. Keep one clock at location $A$ and send the other clock to a location $B$. Suppose that the dilaton field has different values $\chi_{A}$ and $\chi_{B}$ at the two locations $A$ and $B$, respectively. At their respective locations, the clocks would run at different rates given by
$\tau_{A}^{-1}\propto\chi_{A}^{3/2};\ \ \ \tau_{B}^{-1}\propto\chi_{B}^{3/2}\,.$ (30)
When the clock from location $B$ is brought back to location $A$, it will have registered a different amount of elapsed time compared to the clock that resides at location $A$ throughout the experiment, thereby resulting in a new time dilation effect.
This predicted effect differs from the time dilation effect in GR, which is associated with the $g_{00}$ component of the spacetime metric [18, 19]. The new effect arises from the dependence of the clock rate on the dilaton field $\chi$, per Eq. (30). We investigate this new distinct phenomenon further elsewhere.  [^3] In addition to the time dilation effect in GR, there is a well-known effect in Special Relativity where two twice-intersecting time-like paths can have different total amounts of proper time in between. Therefore, we refer to our predicted time dilation effect as “the Third kind”.
Revisiting the Fujii–Wetterich scheme—We must emphasize that the dependence of clock rates on $\chi$ has been documented in the works of Fujii and Wetterich [4, 5, 6, 7, 8, 9], although it was not a focal point in their analysis. However, their findings significantly diverge from ours. Specifically, they reported the following relation for the clock rate  [^4] It is illuminating to quote Fujii [4]: “… the time and length in the microscopic unit frame are measured in units of $m^{-1}(t)$, in agreement with the physical situation that the time scale of atomic clocks, for example, is provided by the atomic levels which are determined by the Rydberg constant $(me^{4})^{-1}$ and Wetterich [7]: “The clock provided by the Hubble expansion in the standard description is now replaced by a clock associated to the increasing value of $\chi$.
$\tau_{\,\text{FW}}^{-1}\propto\chi$ (31)
This contrasts with our result $\tau^{-1}\propto\chi^{3/2}$, per Eq. (28).
Their result can be deduced as follows. If our analysis is repeated using the FW scheme (which mandates that $\hbar$ and $c$ be constant), utilizing Eq. (11), one would obtain for the Bohr radius
$a_{B}=\frac{\hbar}{\alpha m_{\chi}c}=\frac{1}{\alpha\mu\chi}$ (32)
and for the energy level
$E_{n}^{j}=\mathcal{N}_{n}^{j}\,m_{\chi}c^{2}=\mathcal{N}_{n}^{j}\,\mu\hbar c\,\chi\,.$ (33)
These expressions are identical to our results as given in Eqs. (21) and (23), recalling that $\hbar$ and $c$ are constants for the FW scheme. However, the frequency of the photon emitted during the transition of a hydrogen atom from the excited state $\left|n=2,\,j=3/2\right\rangle$ to the groundstate $\left|n=1,\,j=1/2\right\rangle$ is
$\nu_{\,\text{FW}}=\frac{\Delta E}{\hbar}=\Bigl{(}\mathcal{N}_{n=2}^{j=3/2}-\mathcal{N}_{n=1}^{j=1/2}\Bigr{)}\,\mu c\,\chi$ (34)
which results in $\tau_{\,\text{FW}}:=\nu_{\,\text{FW}}^{-1}\propto\chi^{-1}$, the result stated in Eq. (31).
Therefore, despite starting from the same matter Lagrangian $\mathcal{L}_{\text{QED}}^{(\chi)}$, our scheme and the FW scheme are not physically equivalent. They produce two decisively different predictions regarding the behavior of clock rates. Future technologies may be able to distinguish the two predictions and determine the validity of each scheme.
Advantages of our scheme—There are three distinct benefits that our scheme offers as compared to the FW scheme:
1. Equal treatment of mass and charge: The FW scheme treats particle masses and charges at disparity: while masses are promoted to scalar fields proportional to $\chi$, charges are treated as parameters (per Eq. (11)). In contrast, our scheme considers both inertial mass and gauge charge—intrinsic properties of a particle—on equal footing as parameters rather than fields.
2. Equal treatment of $G$ and $e$: In the FW scheme, $G\propto\chi^{-2}$; however, in our scheme, $G$ is a parameter on equal footing with the gauge charge $e$. Consequently, our scheme suggests a commonality between gravitational and gauge interactions.
3. Universal impact on all particle types: The FW scheme applies only to massive particles, leaving massless particles unaffected. In contrast, our scheme allows the dilaton—through its role in determining $\hbar$ and $c$—to influence both massive and massless particles equally.
Discussion—Scalar degrees of freedom naturally emerge in various theoretical frameworks, including Kaluza–Klein, string theory, and braneworld scenarios [20]. By allowing matter to couple non-minimally with gravity via a scalar field (e.g., $\chi$), scalar–tensor theories can acquire scale symmetry, enabling them to evade observational constraints related to the fifth force [2, 3]. On the ground of dimensionality, it can be deduced that the lengthscale $l$ of a given physical process is determined by the value of $\chi$, per $l\propto\chi^{-1}$. This relationship justifies the use of the term “dilaton” for the field $\chi$.
In this paper, we imposed two requirements: (i) A dilaton field $\chi$ exists and directly couples with matter; and (ii) The intrinsic properties of matter—i.e., its gauge charge and inertial mass—are independent of the dilaton. Using the Lagrangian $\mathcal{L}_{\text{QED}}^{(\chi)}$ as a prototype, we demonstrated that these two prerequisites unambiguously lead to the following dependencies of $\hbar$ and $c$ on $\chi$: $\hbar_{\chi}\propto\chi^{-1/2}$ and $c_{\chi}\propto\chi^{1/2}$.
Based on these relationships, we then established that the timescale $\tau$ of a given physical process is related to the value of $\chi$ as $\tau\propto\chi^{-3/2}$. Together with $l\propto\chi^{-1}$, this leads to a universal anisotropic scaling law $\tau\propto l^{3/2}$ between the timescale and lengthscale of a physical process, resulting in a new time dilation effect. (Note: notwithstanding the resemblance of our anisotropic timescaling law to that in the Hořava–Lifshitz gravity [21], Lorentz symmetry is not broken in our approach.)
Three remarks are in order. Firstly, although $h_{\chi}$ and $c_{\chi}$ vary alongside $\chi$ in spacetime, the product $\hbar_{\chi}c_{\chi}$ remains constant. Secondly, in place of the QED prototype considered in this paper, an extension that incorporates the Higgs field and spontaneous symmetry breaking was presented in Ref. [22]. Its generalization to cover the Glashow–Weinberg–Salam model operating under electroweak symmetry breaking is also straightforward.
Finally, and most importantly, the stated dependencies of $\hbar$ and $c$ on $\chi$ are derived solely from the matter Lagrangian without requiring detailed knowledge of the dynamics of $\chi$. This feature renders our derivation particularly powerful and versatile:  the key requirement is the existence of a dilaton that directly couples with matter, whereas the gravitational Lagrangian $\mathcal{L_{\chi}}$, which governs the dynamics of the dilaton, does not play any essential role in our derivation. This versatility opens the door to exploring other candidates for the gravitational Lagrangian in future research, all while maintaining the ubiquity of variable $\hbar$ and $c$.
Physical consequences—A varying $c$ in spacetime leads to direct and significant ramifications in theoretical and observational cosmology. Specifically, a variation in $c$ affects the propagation of light rays in an expanding universe, thereby altering standard cosmography both qualitatively and quantitatively. In [22, 23], we consider a VSL cosmology that relates the speed of light to the cosmic scale factor as $c\propto a^{-1/2}$. For this cosmology, we find that the canonical Lemaître redshift formula,$1+z=a^{-1}$, is no longer applicable and should be modified to $1+z=a^{-3/2}$. The $3/2$–exponent in this new formula arises from the anisotropic time scaling discussed in this paper. Consequently, our new Lemaître redshift formula necessitates a reanalysis for the Hubble diagram of Type Ia supernovae in the context of late-time accelerating expansion. In [22, 23], we conduct this reanalysis for the Pantheon Catalog using our VSL cosmology instead of the concordance $\Lambda$CDM model, and provide a new robust explanation for late-time cosmic acceleration based on VSL, while bypassing the need for dark energy.
Additionally, a varying $\hbar$ may have significant implications for quantum fields in curved spacetimes. One promising avenue for future research on this front is the thermodynamics of black holes in scale-invariant gravity.
Acknowledgements.
I thank Cliff Burgess, Tiberiu Harko, Robert Mann, and Anne-Christine Davis for their constructive and supportive comments during the development of this work.